Discrete Mathematics & Theoretical Computer Science Copyright Transfer

Author(s):

Title of the paper:

I hereby transfer to the Discrete Mathematics and Theoretical Computer Science, DMTCS, non profit organization constitued under French law, all the rights concerning the paper herein mentionned. I confirm, that my co-authors^{1} and their employer(s)^{1} agreed on this step.

I am aware of the fact that the DMTCS guarantees free access to the paper on the internet and that the copyright falls back to the authors in case the internet service of DMTCS can not be continued.

I am also aware, that the authors, as anybody else, for non-profit use only have the right to copy the electronic version of the paper as it is on the server of DMTCS and to distribute it freely in its electronic form.

Date: Address:

Signature:

Discrete Mathematics and Theoretical Computer Science (DMTCS)

62, rue du Cardinal Mathieu F-54000 Nancy – France

1Please rule out if not applicable.